# appendixapcList of Appendices

### v2

@@ 4.2 Human and LLM surprise cluster similar tokens together
Kumar et al. [2023] found strong correlations between human-perceived events and prediction errors across 3 short podcasts (7-30 minutes on average), when processing the corresponding transcript with an LLM. Taking advantage of such human-annotated data and results from previous studies on this dataset [Michelmann et al., 2021, Lositsky et al., 2016], we compare the segmentation quality and correlation with human segmentation for each of our methods (4) using our similarity metrics.
Figure 4: Comparison of human event segmentation with different computational segmentation methods in a human-annotated audio dataset (see also Appendix B). (A) Difference in metrics for the cohesion and separation of KV cache of each LLaMA2 layer. The graphs report the difference of each method with the corresponding random segmentation. (B) Distance between human reports and different methods. In both sets of results, fixed methods (F, FM, FC) perform worse than their surprise-based counterparts (S, SM, SC) with InfLLM’s method (F) performing worse than random.
@@ 4.4 Similarity, Contiguity, Recency and Temporal Order
Table 2: Comparison with graph-theoretic metrics in the KV cache of different LLMs and segmentation methods using the PG-19 dataset and $\gamma=10^{-3}$. Reported values are the difference with random segmentation. Mod: modularity×$\times 10^{5}$, Con: conductance, I/IS: intra/inter-similarity ×$\times 10^{3}$.
@@ 5 Discussion
Significant correlations have been found between human event segmentation and prediction errors in both LLMs [Kumar et al., 2023] and video models [Fountas et al., 2022, Mariola et al., 2022]. Our results add to this growing body of evidence, demonstrating that LLM-based surprise can serve as a proxy for human event segmentation, in multiple levels of hierarchical abstraction, and that the resulting event structure in EM-LLM’s attention heads correlates strongly with human-perceived events. This finding creates a more direct, low-level connection between LLM mechanisms and human cognitive processes. Furthermore, our model’s use of both similarity-based and temporally contiguous retrieval mechanisms parallels human memory retrieval patterns, allowing for the expression of robust phenomena found in human memory research [Howard and Kahana, 2002]. The temporal contiguity effect, where items experienced close together in time are often recalled together, is a robust phenomenon in human memory research [Howard and Kahana, 2002]. Further experiments could deepen our understanding of the connections between EM-LLM and human episodic memory. Following Michelmann et al. [2023b], one could test whether the timing of the event boundaries or the degree of modularity per level that our method produces is closer on average to the human consensus, than individual human subjects. Additionally, exploring how different ratios of the contiguity buffer affect the reproduction of human memory biases, and investigating the impact of recency and initial surprise on event recall, could reveal the extent to which EM-LLM exhibits biases found in free recall studies.
@@ A.1 Comparison with KV-retrieval-based LLMs
Task InfLLM Max Imp. EM-LLM
@@
NarrativeQA 22.12 1.49% 21.77 21.13 21.10 22.45
Qasper 29.33 0.17% 29.07 29.38 29.16 28.68
MultiFieldQA 47.42 1.46% 48.11 47.39 47.72 47.62
HotpotQA 36.56 10.15% 39.40 39.01 40.27 38.90
2WikiMQA 22.31 5.20% 23.46 22.75 23.47 23.46
Musique 17.68 6.17% 17.97 17.82 17.98 18.77
GovReport 31.03 1.90% 31.40 31.62 31.10 31.43
QMSum 23.49 2.13% 23.99 23.20 23.48 23.47
MultiNews 26.70 -0.30% 26.55 26.54 26.58 26.62
TREC 69.00 2.90% 71.00 70.00 70.50 70.50
TriviaQA 86.67 1.10% 86.58 87.62 87.52 87.47
SAMSum 42.52 0.45% 42.71 42.13 42.34 42.48
PassageRetrieval 64.00 32.69% 82.67 78.92 84.92 84.08
LCC 56.67 0.64% 55.03 57.03 54.90 56.79
RepoBench-P 52.97 1.34% 50.49 53.68 51.06 52.86
@@
Code.Debug 29.44 0.88% 29.70 28.43 28.68 28.17
Math.Find 26.57 5.38% 28.00 27.43 27.71 27.14
Retrieve.KV 95.60 3.56% 92.20 97.20 97.60 99.00
En.MC 43.23 -1.02% 41.48 40.61 42.79 42.79
Retrieve.PassKey 100.00 0.00% 100.00 100.00 100.00 100.00
Retrieve.Number 99.83 0.00% 99.83 99.83 99.83 99.83
@@ A.2 Comparison with RAG
In our experiments with Retrieval-Augmented Generation (RAG) baselines, we implemented a standard RAG pipeline consisting of a retriever and a downstream LLM. For each example in a benchmark task, the example context is split into chunks of $l$ words each and encoded using the retriever’s embedding model into a vector database. A similarity lookup into the vector database is used to retrieve the top $k$ most relevant chunks, which are then fed to the downstream LLM alongside the query and task description. For all experiments, we set $l=300$ and $k=5$, following the protocol of Li et al. [2024c].
@@
Task RAG FC EM-LLM
NarrativeQA 22.54 29.14 26.05
Qasper 45.45 45.34 44.41
MultiFieldQA 51.67 54.98 52.52
HotpotQA 55.93 54.01 54.02
2WikiMQA 42.93 45.95 45.72
Musique 30.90 33.52 25.37
GovReport 29.91 34.49 35.04
QMSum 24.97 25.14 24.31
MultiNews 26.77 27.00 27.76
TREC 22.50 4.50 71.50
TriviaQA 88.11 89.07 92.34
SAMSum 7.56 8.68 43.31
PassageRetrieval 65.50 100.00 99.50
LCC 13.16 19.30 67.45
RepoBench-P 18.66 18.33 64.33
Avg. score: 36.44 39.30 51.58
Code.Debug 22.59 21.70 22.59
Math.Find 35.43 26.29 36.00
Retrieve.KV 31.80 92.60 96.80
En.MC 64.19 58.07 44.54
Retrieve.PassKey 100.00 100.00 100.00
Retrieve.Number 99.83 99.32 100.00
Avg. score: 58.97 66.33 66.66
Table 9: EM-LLMS (4K+4K) vs. RAG (NV-Embed-v2 retriever) vs. full-context, with LLaMa-3.1-8B as the base LLM, evaluated on LongBench and $\infty$-Bench.
@@ Appendix C Complexity
C.1 Complexity analysis
@@
Here, we provide a detailed analysis of the computational complexity of our Algorithm 1, focusing on the boundary refinement step and the calculation of modularity and conductance metrics.
Boundary Refinement Step
@@
The boundary refinement step involves finding the optimal position $\hat{\beta}$ between each pair of consecutive initial boundaries $(\alpha,\beta)$ that optimizes the chosen metric function $f$. This step has the following components:
Iteration over initial boundaries: $\mathcal{O}(k)$, where $k$ is the number of initial boundaries. For each pair of boundaries, we compute the metric function $f$ for all positions between $\alpha$ and $\beta$. In the worst case, this is $\mathcal{O}(n)$ operations per boundary pair.
Therefore, the overall complexity of this step is $\mathcal{O}(kn)$.
Metric Function Computation
The metric functions (modularity or conductance) are computed at the level of individual memory units. For a memory unit of size $m$:
  • Modularity: The naive computation involves summing over all pairs of nodes within the unit, resulting in a worst-case complexity of $\mathcal{O}(m^{2})$. However, in practice, the similarity graph is often sparse, meaning many node pairs have negligible similarity. Leveraging this sparsity, more efficient implementations can achieve $\mathcal{O}(l)$ complexity, where $l$ is the number of non-zero similarity edges within the unit [Newman, 2004]. Typically, $l$ is much smaller than $m^{2}$, especially for larger units, leading to significant computational savings.
  • Conductance: This requires computing the sum of edge weights crossing the boundary and the total volume of the unit, which can be done in $\mathcal{O}(m)$ time.
Given that $m$ is typically much smaller than $n$ and varies for each unit, we can consider the average unit size $\bar{m}$ and average number of non-zero similarity edges $\bar{l}$. The total complexity for computing metrics across all units is then $\mathcal{O}(k\bar{l})$ for modularity (which in the worst case is $\mathcal{O}(k\bar{m}^{2})$, but typically much lower) or $\mathcal{O}(k\bar{m})$ for conductance.
Overall Complexity
Combining the boundary refinement step and metric computation, the overall complexity is:
For modularity: $\mathcal{O}(kn+k\bar{m}^{2})$ For conductance: $\mathcal{O}(kn+k\bar{m})$
Since typically $\bar{m}\ll n$, the dominant term in both cases is $\mathcal{O}(kn)$. Therefore, we express the overall complexity of our algorithm as $\mathcal{O}(kn)$.
C.2 Implementation and use of hardware resources
@@
C.2.1 Wall Clock Time
@@
Table 10: Difference in wall-clock time to process a single chunk of 512 tokens for various ablations of our framework as compared to InfLLM. Measured using Mistral-7B (4K+2K) and averaged over the first 100 chunks ($51.2$K tokens) of a long sequence.
As can be observed in Table 10, the largest increase to wall clock time in our framework is due to the similarity adjustment step. We would also like to note that such times increase steadily as sequence length increases, due to the increasing number of representative tokens involved in the k-NN calculation used for block retrieval, regardless of which method is used.
C.2.2 Memory Requirements
@@ Appendix D Further Ablations
D.1 Surprise, refinement and contiguity

### v3

@@ 4.2 Human and LLM surprise cluster similar tokens together
Kumar et al. [2023] found strong correlations between human-perceived events and prediction errors across 3 short podcasts (7-30 minutes), when processing the corresponding transcript with an LLM. Taking advantage of such human-annotated data and results from previous studies on this dataset [Michelmann et al., 2021, Lositsky et al., 2016], we compare the segmentation quality and correlation with human segmentation for each of our methods (Fig. 4) using our similarity metrics.
Figure 4: Comparison of human event segmentation with different computational segmentation methods in a human-annotated audio dataset (see also Appendix B). (A) Difference in metrics for the cohesion and separation of KV cache of each LLaMA2 layer. The graphs report the difference of each method with the corresponding random segmentation. (B) Distance between human reports and different methods. In both sets of results, fixed methods (F, FM, FC | with M: Modularity, C: Conductance) perform worse than their surprise-based counterparts (S, SM, SC) with InfLLM’s method (F) performing worse than random.
@@ 4.4 Similarity, Contiguity, Recency and Temporal Order
[…]
@@ 5 Discussion
Significant correlations have been found between human event segmentation and prediction errors in both LLMs [Kumar et al., 2023] and video models [Fountas et al., 2022, Mariola et al., 2022]. Our results add to this growing body of evidence, demonstrating that LLM-based surprise can serve as a proxy for human event segmentation, in multiple levels of hierarchical abstraction, and that the resulting event structure in EM-LLM’s attention heads correlates strongly with human-perceived events. This finding suggests a potential, low-level parallels between LLM mechanisms and human cognitive processes (see also Appendix E.1). Furthermore, our model’s use of both similarity-based and temporally contiguous retrieval mechanisms parallels human memory retrieval patterns, allowing for the expression of robust phenomena found in human memory research [Howard and Kahana, 2002]. The temporal contiguity effect, where items experienced close together in time are often recalled together, is a robust phenomenon in human memory research [Howard and Kahana, 2002]. Further experiments could deepen our understanding of the connections between EM-LLM and human episodic memory. Following Michelmann et al. [2023b], one could test whether the timing of the event boundaries or the degree of modularity per level that our method produces is closer on average to the human consensus, than individual human subjects. Additionally, exploring how different ratios of the contiguity buffer affect the reproduction of human memory biases, and investigating the impact of recency and initial surprise on event recall, could reveal the extent to which EM-LLM exhibits biases found in free recall studies.
@@ A.1 Comparison with KV-retrieval-based LLMs
While the vast majority of our results in this section show a statistically significant improvement on InfLLM at the benchmark level ($p<0.05$ using a two-tailed z-test, except LongBench with Phi-3.5 with $p=0.23$), it should be noted this isn’t the case in the majority of individual tasks. However, given the consistency and frequency of improvements across a large number of such tasks, along with the benchmark-level significance of such improvements, we consider the lower, task-level significance to be largely due to the sample size of individual tasks rather than chance, and believe it is still reasonable and justified to claim an overall improvement on InfLLM. Moreover, including individual task results supports transparency and allows for future works to make more granular comparisons and use of such results.
Task Type Task InfLLM Max Imp. EM-LLM
@@
Single-doc QA NarrativeQA 22.12 1.49% 21.77 21.13 21.10 22.45
" Qasper 29.33 0.17% 29.07 29.38 29.16 28.68
" MultiFieldQA 47.42 1.46% 48.11 47.39 47.72 47.62
Multi-doc QA HotpotQA 36.56 10.15% 39.40 39.01 40.27 38.90
" 2WikiMQA 22.31 5.20% 23.46 22.75 23.47 23.46
" Musique 17.68 6.17% 17.97 17.82 17.98 18.77
Summarisation GovReport 31.03 1.90% 31.40 31.62 31.10 31.43
" QMSum 23.49 2.13% 23.99 23.20 23.48 23.47
" MultiNews 26.70 -0.30% 26.55 26.54 26.58 26.62
Few shot TREC 69.00 2.90% 71.00 70.00 70.50 70.50
" TriviaQA 86.67 1.10% 86.58 87.62 87.52 87.47
" SAMSum 42.52 0.45% 42.71 42.13 42.34 42.48
Retrieval PassageRetrieval 64.00 32.69% 82.67 78.92 84.92 84.08
Code LCC 56.67 0.64% 55.03 57.03 54.90 56.79
" RepoBench-P 52.97 1.34% 50.49 53.68 51.06 52.86
@@
Code Code.Debug 29.44 0.88% 29.70 28.43 28.68 28.17
Multiple choice En.MC 43.23 -1.02% 41.48 40.61 42.79 42.79
Retrieval Math.Find 26.57 5.38% 28.00 27.43 27.71 27.14
" Retrieve.KV 95.60 3.56% 92.20 97.20 97.60 99.00
" Retrieve.PassKey 100.00 0.00% 100.00 100.0 100.00 100.00
" Retrieve.Number 99.83 0.00% 99.83 99.83 99.83 99.83
@@ A.2 Comparison with RAG
In our experiments with Retrieval-Augmented Generation (RAG) baselines, we implemented a standard RAG pipeline consisting of a retriever and a downstream LLM. For each example in a benchmark task, the example context is split into chunks of words each and encoded using the retriever’s embedding model into a vector database. A similarity lookup into the vector database is used to retrieve the top $k$ most relevant chunks, which are then fed to the downstream LLM alongside the query and task description.
@@
For most experiments, we set chunk size to $l=300$ and $k=5$, following the protocol of Li et al. [2024c]. As a comparative baseline, we also trialled chunking according to surprise, the results of which are shown in Table 9 (RAG-S). In this experiment, the context was first segmented into blocks by the EM-LLM (S) model, each of which then encoded by the retriever’s embedding model. Context was dynamically retrieved to fill a buffer of 1500 words, for fair comparison with fixed-size chunking. The RAG-S variant underperformed the fixed-size implementation - we hypothesise that this was due to the retrieved context in RAG-S consisting of highly disjointed information, due to small chunk sizes and high $k$. Incorporating contiguity into the retrieval mechanism may be sufficient to close this performance gap.
@@
Task RAG-S RAG FC EM-LLM
NarrativeQA 12.10 22.54 29.14 26.05
Qasper 36.41 45.45 45.34 44.41
MultiFieldQA 44.08 51.67 54.98 52.52
HotpotQA 41.56 55.93 54.01 54.02
2WikiMQA 28.84 42.93 45.95 45.72
Musique 19.04 30.90 33.52 25.37
GovReport 18.12 29.91 34.49 35.04
QMSum 19.22 24.97 25.14 24.31
MultiNews 26.21 26.77 27.00 27.76
TREC 2.5 22.50 4.50 71.50
TriviaQA 88.26 88.11 89.07 92.34
SAMSum 8.09 7.56 8.68 43.31
PassageRetrieval 16.0 65.50 100.00 99.50
LCC 11.02 13.16 19.30 67.45
RepoBench-P 17.39 18.66 18.33 64.33
Avg. score: 25.89 36.44 39.30 51.58
Code.Debug - 22.59 21.70 22.59
Math.Find - 35.43 26.29 36.00
Retrieve.KV - 31.80 92.60 96.80
En.MC - 64.19 58.07 44.54
Retrieve.PassKey - 100.00 100.00 100.00
Retrieve.Number - 99.83 99.32 100.00
Avg. score: - 58.97 66.33 66.66
Table 9: EM-LLMS (4K+4K) vs. RAG (NV-Embed-v2 retriever) vs. full-context, with LLaMa-3.1-8B as the base LLM, evaluated on LongBench and $\infty$-Bench. The comparison also includes RAG-S, which is the same RAG retriever but with the same surprise-based segmentation used in EM-LLM results.
@@ Appendix C Complexity
C.1 Boundary Refinement Complexity Analysis
@@
Here, we provide a detailed analysis of the computational complexity of our Algorithm 1, focusing on the boundary refinement step and the calculation of modularity and conductance metrics. Here we describe scaling complexity with example context length $n$, a chunk size $m$, and $k$ events.
Metric Function Computation The metric functions (modularity or conductance) are computed at the level of individual memory units but all rely on the same adjacency matrix for that chunk. The complexity of calculating the adjacency matrix, as defined in Eq. 2, is $\mathcal{O}(m^{2})$. As the complexity for both metrics is largely dominated by, or on the same order as the computation of this term, the resulting complexity for the calculation of the metric function is $\mathcal{O}(m^{2})$.
@@
Boundary Refinement Step The boundary refinement step involves finding the optimal position $\hat{\beta}$ between each pair of consecutive initial boundaries $(\alpha,\beta)$ that optimizes the chosen metric function $f$. The iteration over initial boundaries has complexity $\mathcal{O}(k)$. For each updated boundary we change the community of one node at at time to the one on its right and re-evaluate the metric function $f$ with this change. Hence, for each boundary position, on average $2\frac{n}{k}$ nodes see a change in their community, and hence need to re-evaluate their contribution to the metric function. This results in a complexity of $\mathcal{O}(\frac{n}{k})$, for each position between $\alpha$ and $\beta$. This step will therefore scale with average event size resulting in a complexity $\mathcal{O}((\frac{n}{k})^{2})$. Therefore, the overall complexity of this step scales with $\mathcal{O}(k(\frac{n}{k})^{2})=\mathcal{O}(\frac{n^{2}}{k})$.
Overall Complexity The context is divided into $\frac{n}{m}$ chunks, and hence, we must compute this number of initial adjacency matrices with resulting complexity $\mathcal{O}(\frac{n}{m}m^{2})=\mathcal{O}(nm)$. Adding in the complexity of the refinement updates, we get $\mathcal{O}(nm+\frac{n^{2}}{k})$. In practice, our method detects $k$ events such that $\frac{n}{k}\leq m$ is always true, hence this is upper-bounded by $\mathcal{O}(nm)$.
C.2 Attention Complexity Analysis
Attention. Calculating a full attention matrix for a sequence length $n$ has complexity $\mathcal{O}(n^{2})$. However, in the case of EM-LLM, we process such a context in smaller chunks of size $m$ and rely on retrieval to approximate a full attention matrix (also see Appendix F.1). Therefore we evaluate $\frac{n}{m}$ attention matrices each with complexity $\mathcal{O}(m(n_{l}+n_{r}))$ with $n_{l}$ the number of local tokens used and $n_{r}$ the number of tokens retrieved as part of the multi-stage attention process.
k-NN Retrieval. As previously mentioned, our approach relies on k-NN retrieval to avoid computing a full attention matrix. Such a retrieval process scales with the maximum number of memory units saved. In our case, as we enforce a minimum block size $b$, the maximum possible number of memory units to check for retrieval is $\frac{n-n_{l}}{b}$. As we retrieve memory units for every processed chunk (past $n_{l}$ processed tokens) this is upper-bounded by $\mathcal{O}(\frac{n(n-n_{l})}{mb})$.
Overall. Including the retrieval step, overall our approach performs its attention computation with complexity $\mathcal{O}(n(n_{l}+n_{r})+\frac{n(n-n_{l})}{mb})$. In practice, for long-context tasks, this scales much slower than a full attention matrix, as visualized in Figures 9 and 10 which demonstrate this with our own settings and for values of $n$ up to 100K and 10M respectively. In fact, our complexity is negligible compared to full attention once sequence length reaches the millions.
Figure 9: A visualization of the scaling complexity of EM-LLM vs a standard full-context approach (full attention matrix) as a function of sequence length (up to 100K tokens).
Figure 10: A visualization of the scaling complexity of EM-LLM vs a standard full-context approach (full attention matrix) as a function of sequence length (up to 10M tokens).
C.3 Implementation and use of hardware resources
@@
C.3.1 Wall Clock Time
@@
Table 10: Difference in wall-clock time to process a single chunk of 512 tokens for various ablations of our framework as compared to InfLLM. Measured using Mistral-7B (4K+2K) and averaged over the first 100 chunks (51.251.2K tokens) of a long sequence.
As can be observed in Table 10, the largest increase to wall clock time in our framework is due to the similarity adjustment step. We would also like to note that such times increase steadily as sequence length increases, due to the increasing number of representative tokens involved in the k-NN calculation used for block retrieval, regardless of which method is used, although this is only noticeable when sequence length reaches the millions (see Appendix C.2).
C.3.2 Memory Requirements
@@ Appendix D Further Ablations
D.1 Hyper-parameter Selection and Tuning
Surprise Threshold Parameter $\gamma$
Equation 1 introduces the surprise threshold parameter $\gamma$, which is responsible for the sensitivity of the threshold by scaling the standard deviation measured across the moving window. As such, with $\gamma=1$, for a new token do be considered "surprising" it must achieve a surprise value greater than one standard deviation over the moving average. As mentioned in section 3.2, this ensures that the threshold adapts to contextual shifts therefore minimizing the need for manual tuning.
We initially explored our approach’s sensitivity to $\gamma$ using Mistral on the LongBench benchmark. We evaluated the benchmark using surprise-only segmentation with $\gamma\in{1.0,1.5,2.0,2.5,3.0,3.5}$, as visualized in Figure 12. Naturally, we noticed that an increase in $\gamma$ resulted in a decrease in the number of events detected, and a resulting increase in the mean event size. In terms of overall performance, results suggest that a smaller $\gamma$, is generally best. We then explored such behavior across our segmentation methods, for which the overall behaviors were largely consistent with the surprise-only method. One particularly interesting observation is the fact the addition of our refinement algorithm ($\mathbf{sM}$) seemed to show particularly high improvements over surprise-only at larger values of $\gamma$, although the best-performing value was still $\gamma=1$.
Following these initial observations, in order to choose $\gamma$ for our experiments, we evaluated each model on the LongBench benchmark with $\gamma\in{1,2,3}$ and surprise-only segmentation. We then selected the best-performing value of $\gamma$ for the rest of our experiments with each model. These were $\gamma=1,2,1,1,1$ for Mistral, LLaMa-3, LLaMa-3.1, Phi-3, Phi-3.5 respectively, showing a consistent preference for $\gamma=1$.
Retrieved vs. Contiguity Buffer Ratio
Section 3.4 introduces a similarity buffer of $k_{s}$ events and a contiguity buffer of $k_{c}$ events which retrieves $n$ events either side of each event in the similarity buffer. In total, the number of retrieved events is $k=k_{s}+k_{c}$. Note that we chose $n=1$ for all experiments in order to balance contiguity with similarity. Moreover, Fig. 3A&B suggest that the contiguity effect is most significant for a handful of positions either side of the one recalled. This suggests that larger values of $n$ are likely to bring diminishing returns in terms of contiguity, while also reducing the size of the similarity buffer.
In practice, $k$ is expressed as a maximum number of tokens to ensure consistency as event sizes vary due to our dynamic segmentation methods. We therefore express the sizes of the similarity and contiguity buffers as a ratio $k_{r}=\frac{k_{c}}{k}$, and dynamically fill them with events such that $k_{s}=(1-k_{r})k$ and $k_{c}=k_{r}k$. In order to find the best contiguity ratio, we evaluated LongBench with Mistral and $\gamma\in{1,2}$ over values $k_{r}\in{0.3,0.5,0.7}$, as illustrated in Fig. 13. To give a bit of intuition as to the meaning of these values, $k_{r}=0.3$ will, on average, correspond to only including contiguous events for the top $25\%$ of events in the similarity buffer. On the other hand, $k_{r}=0.7$, on average, will correspond to including contiguous events for all events in the similarity buffer, but will include $50\%$ less events in the latter. As mentioned in section 4.4, contiguity appears to have varying importance across tasks, with some performing best with the larger $k_{r}=0.7$, but most did best with the lower $k_{r}=0.3,0.5$. Results also showed a clear preference for $\gamma=1$, which is consistent with our previous experiments. Overall, there was a slight preference for $k_{r}=0.3$, which we therefore selected for the rest of our experiments.
D.2 Surprise, refinement and contiguity